# Generalized Reduced-Density-Matrix Quantum Monte Carlo Gives Access to More

### 11 Mar 2026 v1

@@ Abstract
In quantum Monte Carlo (QMC), what can be measured efficiently is largely determined by what is sampled. When the sampled object is the partition function, a broad class of observables, including general off-diagonal operators, is typically unavailable as direct estimators. In this article, we introduce a paradigm shift by replacing the partition function with a generalized reduced density matrix (GRDM) as the simulated object. This reformulation removes the measurement bottleneck at its source and extends the dimensional-reduction advantage of reduced descriptions from static quantities to dynamical observables, thereby enabling much richer information extraction. As substantial demonstrations, the framework allows the directed-loop algorithm to measure both equal-time and imaginary-time off-diagonal observables, with the latter giving direct access to dynamical spectra. It also enables measurements of Rényi-1 correlators that diagnose strong-to-weak symmetry breaking in mixed states. This work establishes a unified framework for holographic characterization within QMC.
@@ Appendix C Appendix C: GRDM method and off-diagonal correction function
Furthermore, this framework is not restricted to imaginary-time correlations. Equal-time observables are obtained simultaneously. By isolating and treating the $\rho_{\hat{I}}$ component from Eq. (16) independently, we obtain:

### 12 Mar 2026 v2

@@ Abstract
For a long time, people have been focusing on how to extract more information, such as off-diagonal observables, from the quantum Monte Carlo (QMC) simulation of the partition function, but there have been numerous difficulties, and many of them are insurmountable. In this article, we point out that all the difficulties stem from the starting point of the simulation: calculating a partition function. We introduce a paradigm shift: when we transform the simulated object from a partition function to a generalized reduced density matrix (GRDM), the difficult problem of measurement can be readily solved. By designing the GRDM, both equal-time and nonequal-time off-diagonal observables have been measured easily in QMC with a polynomial computation complexity. As a demonstration, the GRDM enables direct access to nonequal-time correlators for dynamical spectra as well as Rényi-1 correlators that reveal strong-to-weak symmetry breaking in the mixed state, capabilities that lie beyond the reach of prior methods. This establishes a unified framework for holographic characterization within QMC.
@@ Appendix C Appendix C: GRDM method and off-diagonal correction function
Furthermore, this framework is not restricted to imaginary-time correlations. Equal-time observables are obtained simultaneously. By isolating and treating the $\tilde{\rho}_{\hat{I}}$ component from Eq. (16) independently, we obtain: